# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v2

@@ 3.2 Bounded-Genus Graphs
Kelner’s bound on $\lambda_{2}$ (stated in Theorem 2.3) for triangulated genus-$g$ graphs also makes use of circle packings [Kel04, Kel06b, Kel06a], and, as we will see, can also be generalized to a bound on $\lambda_{2}^{*}$ and $\gamma^{(1)}$. However, the resulting bound also has a (large) polynomial dependence on the maximum degree $\Delta$. Rather than directly applying this bound, we will apply a new transformation that allows us to almost fully tighten this dependence. Specifically, our new transformation will take a genus-$g$ graph with $n$ vertices and maximum degree $\Delta$, and produce a graph with $O(\Delta\cdot n)$ vertices and constant maximum degree. We will then relate the values of $\gamma^{(1)}$ for the two graphs. In the case of $\lambda_{2}^{*}$ and $\gamma^{(1)}$, this relationship will result in a greatly improved $O(\log\Delta)$ factor for the upper bound.
@@
It is notable that Kelner’s bound on $\lambda_{2}$ does not apply to non-triangulated genus-$g$ graphs, despite the paper claiming otherwise. The claimed argument is as follows:
@@
  • Apply the bound on triangulated genus-$g$ graphs that Kelner proves in the rest of his paper. Note that this bound relies on a generalized circle packing theorem of He and Schramm [HS93] that applies only to simple triangulations of genus-$g$ surfaces.
@@
  • A maximal genus-$g$ graph is not always a triangulation. In particular, $K_{5}$ is a maximal genus-$1$ graph that is not triangulated in any embedding into the torus. We may prove this by contradiction: Suppose this was not the case, and that it had a triangulated embedding in the torus. Then, by the Euler characteristic, such an embedding must have $5$ faces. We may count the number of edge-face incidences. Since the embedding is triangulated (and hence cellular), each edge is incident to exactly two faces, and there are $10$ edges and hence $2\cdot 10=20$ incidences in total. Since the embedding is triangulated, there are exactly $3$ edges incident to each face, and hence there are $3\cdot 5=15$ incidences in total – a contradiction.
@@
It is worth noting that neither of these issues arises for planar graphs: Maximal planar graphs are always planar triangulations, and there is a known result that triangulates a planar graph while preserving the maximum degree [KB92, KB97]. For genus $g>0$, we will present a method to bypass these issues with carefully chosen Steiner vertices while tightening and generalizing the bound to $\lambda_{2}^{*}$ and $\gamma^{(1)}$.
@@
Let $p:V^{\prime}\to V$ be the function that certifies the depth-$L$ uniform shallow minor. Let $f_{H},y_{H}$ be a solution to $\gamma^{(1)}(H)$. We will construct a solution $f_{G},y_{G}$ to $\gamma^{(1)}(G)$ with random sampling.
@@
Let $G$ be a graph with genus $g$, $n$ vertices, and maximum degree $\Delta$. Then there exists a graph $H$ with $n\Delta$ vertices and maximum degree $4$ with genus $g$ so that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$.
@@
We create a set $V^{\prime}$ that contains $\Delta$ vertices for each vertex in $V(G)$. Let $p:V^{\prime}\to V(G)$ denote the mapping of these vertices. Likewise, for each edge-vertex pair $(e,v)\in E\times V$ with $e$ incident to $v$, map $(e,v)$ to some vertex in $p^{-1}(v)$, in such a manner that no two distinct edge-vertex pairs are mapped to the same vertex in $V^{\prime}$. Use this second mapping to construct a set of endpoint-distinct edges $E^{\prime}$. For each $v\in V(G)$, construct a near-perfect binary tree $T_{v}$ over the $\Delta$ vertices in $p^{-1}(v)$ so that some Euler tour of $T_{v}$ contains the vertices mapped to the rotation system around $v$ as a subsequence (it suffices to construct a near-perfect binary search tree over an arbitrary ordering of $p^{-1}(v)$ containing this subsequence). Note that $T_{v}$ will have max-degree $3$ and diameter at most $2\log_{2}\Delta$. Let $E^{\prime\prime}$ be the union over all such trees. Let $H:=(V^{\prime},E^{\prime}\cup E^{\prime\prime})$. Note that the maximum degree of $H$ is $4$, since each vertex in $V^{\prime}$ is incident to at most one edge in $E^{\prime}$ and at most $3$ edges in $E^{\prime\prime}$.
@@
Figure 3: Two examples of the degree-reduction construction for a genus-$g$ graph, respecting a rotation system, local to a vertex. The vertex is replaced with $\Delta$ vertices, at least one vertex for each incident edge, and these new vertices are connected with a near-perfect binary tree. In the left example, the displayed vertex is of maximum degree in its graph, while in the right example it is not (resulting in some degree-$1$ vertices).
@@
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated genus-$g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $O(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
@@
Let $H_{1}$ be the graph that adds $\Delta$ vertices around each vertex to $G$, and ensure that between any two edges in the rotation system around a vertex in $G$, at least one such edge is added. Observe that $H_{1}$ is a genus-$g$ graph with $(\Delta+1)n$ vertices and maximum degree $2\Delta$. Observe also that $G$ is a uniform shallow minor of $H_{1}$ of depth $2$, and hence Lemma 3.11 shows that $\gamma^{(1)}(G)\lesssim(\Delta+1)\gamma^{(1)}(H_{1})$.
Now, we will add edges to $H_{1}$. In particular, we will construct a sequence of graphs that add edges to $H_{1}$, but never add vertices, and hence $\gamma^{(1)}$ will only increase for this sequence of graphs. Specifically, within each face of $H_{1}$, we create a cycle of all the newly added vertices in order around the face. The result is a graph we call $H_{2}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{3,2\Delta\}$) that has no adjacent pair of faces both with size $>4$. Moreover, any vertex incident to a face of size $>4$ (i.e., those in the created cycles) is incident exactly two other faces, and has degree exactly $3$. For each face of exactly size $4$ in $H_{2}$, triangulate it arbitrarily to get another new graph $H_{3}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{5,4\Delta\}$). See Figure 4 for examples of each of these graphs around a particular face. It’s worth noting that for higher genus graphs, it is possible that a non-simple face may include the same edge twice in the same “direction” during a traversal around the face (as opposed to the example in the figure, where the doubly-included edge is included in opposite directions). This does not cause issues for the construction described here either. It is also worth noting that these steps are necessary even if all the faces are simple: Specifically, we would like to handle the remaining non-triangular faces independently, and the separation ensures that we are not able to later produce parallel edges (see [KB97, Figure 6.3(b)] for a demonstration of the issues of adjacent simple faces).
It remains to triangulate the remaining non-triangular faces of $H_{3}$, none of which are adjacent, and all of which are simple. This can be accomplished by adding at most $2$ edges per vertex. In particular, Kant and Bodlaender [KB92, KB97] gave an algorithm called “zig-zagging” that adds edges to a single face using a sequence of ear-cuts, triangulating the face. This algorithm applies in the case of faces on any rotation system, not just a rotation system on a planar graph. Call the final graph $H$, which is a triangulated genus-$g$ graph with maximum degree $\max\{7,4\Delta\}$. See the last diagram in Figure 4 for a demonstration of this algorithm.
@@
For the remaining lemmas, we need only consider triangulated genus-$g$ graphs.
A genus-$g$ graph naturally embeds into a genus-$g$ surface. Moreover, a genus-$g$ graph in fact has a circle packing on a genus-$g$ surface. Both the surface and packing is unique up to a certain set of transformations [HS93]. However, to bound $\gamma^{(3)}$ in a similar manner to Theorem 3.5, we would need a circle packing on the unit sphere $S^{2}$, likely of ply $O(g)$. Unfortunately, as Kelner [Kel04, Kel06b, Kel06a] pointed out, such circle packings need-not exist. Instead, in order to bound $\lambda_{2}$, Kelner made use of two steps: First, a specific subdivision procedure is applied to the graph, and it is shown that a bound for the graph obtained by a sequence of subdivision procedures also induces a bound on the initial graph. Second, by applying this procedure a sufficient number of times, it is shown that a continuous analogue of the circle packing theory can be used to obtain a bound. We will apply analogues of each of these steps for $\gamma^{(3)}$ (and $\gamma^{(1)}$, which differs only by a constant factor of at most $3$) instead of $\lambda_{2}$.
@@
Let $G$ be a graph of genus $g$. For each $k$, let $F^{(k)}$ denote a genus-$g$ surface on which $G^{(k)}$ admits a circle packing. For each vertex $v\in G^{(k)}$, let $C_{v}$ denote the disk in the packing corresponding to $v$, and let $r_{v}$ denote its radius. Let $A(C)$ and $D(C)$ denote the area (measure) and diameter (longest geodesic path between a point pair), respectively, of a compact and connected region $C$ either in $F^{(k)}$ or $S^{2}$. Then there is a sequence $f^{(k)}:F^{(k)}\to S^{2}$ of analytic maps so that, for any $\varepsilon>0$, there exists a threshold $N$ such that for any $k\geq N$ a partition of $S^{(k)}$ into $S^{(k)}_{1}\cup S^{(k)}_{2}$ exists with the following properties:
@@
Let $G^{\prime}$ be the genus-$g$ graph on $n\Delta$ vertices with maximum degree at most $4=O(1)$ produced by Lemma 3.12. Since the maximum degree of $G^{\prime}$ is constant, a triangulated genus-$g$ graph $G^{\prime\prime}$ with $O(n\Delta)$ vertices and constant maximum degree can be obtained by Lemma 3.13 so that $\gamma^{(1)}(G^{\prime})\lesssim\gamma^{(1)}(G^{\prime\prime})$. By Lemma 3.14, and Lemma 3.16 $\gamma^{(1)}(G^{\prime\prime})\lesssim\frac{g}{n\Delta}$, since the degree of $G^{\prime\prime}$ is constant. Hence,
@@
4 Bounds via $1$-dimensional Extremal $L_{p}$-Embedded Spread
In this section, we will study a quantity we call the $d$-dimensional extremal $L_{2}$-embedded spread (with a focus on the $1$-dimensional case). In more generality:
@@ 4 Bounds via $1$-dimensional Extremal $L_{p}$-Embedded Spread
Apply Theorem 2.15 to get a non-expansive embedding $f^{\prime}:V\to\mathbb{R}$ from $d_{G}^{\omega}$ to $\mathbb{R}$ with average $2$-distortion at most $C_{2}\cdot 2\cdot\left[\overline{\alpha}(G)\right]^{2}$ for a universal constant $C_{2}$. Using $y(v):=\omega(v)^{2}$ and $f(v):=f^{\prime}(v)/2$, we claim the constraints of $\overline{Q}_{2}^{(1)}(G)$ are satisfied: For the first constraint, $||y||_{1}=\sum_{v\in V}\omega(v)^{2}=||\omega||_{2}^{2}\leq 1$. For the remaining constraints,
@@
and so $y,f$ forms a feasible solution to $\overline{Q}_{2}^{(1)}(G)$. Finally, it follows from the $2$-distortion bound that
@@
The proof is essentially the same as that of Lemma 4.5, except that a non-expansive embedding with bounded $1$-distortion is used. ∎
@@ 6 A High-Level View of Congestion-based Bounds
Let $\Lambda^{u,v}$ denote a flow from $u$ to $v$, so that $\Lambda$ itself is a multi-commodity flow for all possible pairs $(u,v)$. $\Lambda$ is said to be a unit $K_{n}$-flow if the amount of flow in $\Lambda^{u,v}$ is exactly $1$ for all pairs $(u,v)$. A unit $K_{n}$-flow $\Lambda$ is said to be integral if all flows are paths. The congestion of a flow $\Lambda$ at a vertex $x$ is $c_{\Lambda}(x):=\sum_{u,v\in V}\Lambda^{u,v}(x)$. The $L^{p}$-extremal congestion is $\text{con}_{p}(G):=\min_{\Lambda}||c_{\Lambda}||_{p}$, where the minimum is taken over all unit $K_{n}$ flows $\Lambda$. The $L^{p}$-extremal integral congestion is $\text{con}_{p}^{*}(G):=\min_{\Lambda}||c_{\Lambda}||_{p}$, where the minimum is taken over all integral unit $K_{n}$ flows $\Lambda$.
@@
$\lambda_{2}^{*}(G)\lesssim(\log g)^{2}\cdot\frac{g}{n}$$\gamma^{(n)}(G)\lesssim(\log g)^{2}\cdot\frac{g}{n}$$\gamma^{(1)}(G)\lesssim(\log g)^{4}\cdot\frac{g}{n}$$\psi(G)\lesssim\frac{(\log g)^{2}\sqrt{g}}{\sqrt{n}}$trong duality Metric embedding/dimension reduction (Theorem 5.2) 11 11-dimensional Cheeger inequality for $\psi$ (Theorem 1.5)
@@ 7 Conclusion and Future Work
Experimentally testing the practicality of the method would first require a complete implementation, so we outline the simplest-known methods for each step: Boyd, Diaconis and Xiao [BDX04] provided a practical and simple subgradient method for computing $\lambda_{2}^{*}$ that could quite easily be implemented with GPU-acceleration in a modern library, although a slightly different method (such as a subgradient method in the dual, or a more recent general-case SDP-solver) would need to be used to allow the extraction of a dual solution for $\gamma^{(n)}$. The simplest dimension-reduction technique is that of Kwok, Lau, and Tung [KLT22], stated in Theorem 1.6. Unfortunately, the best-known bounds for $\alpha(G)$ (particularly for forbidden-minor graphs) are not currently as algorithmically simple, nor is their application in Theorem 5.3. Finally, the simplest cut algorithm is also likely that of Kwok, Lau, and Tung [KLT22, Section 3.2.3].

### v3

@@ 3.2 Bounded-Genus Graphs
Kelner’s bound on $\lambda_{2}$ (stated in Theorem 2.3) for triangulated $g$ graphs also makes use of circle packings [Kel04, Kel06b, Kel06a], and, as we will see, can also be generalized to a bound on $\lambda_{2}^{*}$ and $\gamma^{(1)}$. However, the resulting bound also has a (large) polynomial dependence on the maximum degree $\Delta$. Rather than directly applying this bound, we will apply a new transformation that allows us to almost fully tighten this dependence. Specifically, our new transformation will take a $g$ graph with $n$ vertices and maximum degree $\Delta$, and produce a graph with $O(\Delta\cdot n)$ vertices and constant maximum degree. We will then relate the values of $\gamma^{(1)}$ for the two graphs. In the case of $\lambda_{2}^{*}$ and $\gamma^{(1)}$, this relationship will result in a greatly improved $O(\log\Delta)$ factor for the upper bound.
@@
It is notable that Kelner’s bound on $\lambda_{2}$ does not apply to non-triangulated $g$ graphs, despite the paper claiming otherwise. The claimed argument is as follows:
@@
  • Apply the bound on triangulated $g$ graphs that Kelner proves in the rest of his paper. Note that this bound relies on a generalized circle packing theorem of He and Schramm [HS93] that applies only to simple triangulations of $g$ surfaces.
@@
  • A maximal $g$ graph is not always a triangulation. In particular, $K_{5}$ is a maximal genus-111 graph that is not triangulated in any embedding into the torus. We may prove this by contradiction: Suppose this was not the case, and that it had a triangulated embedding in the torus. Then, by the Euler characteristic, such an embedding must have 555 faces. We may count the number of edge-face incidences. Since the embedding is triangulated (and hence cellular), each edge is incident to exactly two faces, and there are 101010 edges and hence 210=202$\cdot$ 10=202 ⋅ 10 = 20 incidences in total. Since the embedding is triangulated, there are exactly 333 edges incident to each face, and hence there are 35=153$\cdot$ 5=153 ⋅ 5 = 15 incidences in total – a contradiction.
@@
It is worth noting that neither of these issues arise for planar graphs: Maximal planar graphs are always planar triangulations, and there is a known result that triangulates a planar graph while asymptotically preserving the maximum degree [KB92, KB97]. For genus $g>0$, we will present a method to bypass these issues with carefully chosen Steiner vertices while simultaneously tightening and generalizing the bound to $\lambda_{2}^{*}$ and $\gamma^{(1)}$.
@@
[…]
@@
Let $G$ be a graph with genus $g$, $n$ vertices, and maximum degree $\Delta$. Then there exists a graph $H$ with $n\Delta$ vertices and maximum degree 444 with genus $g$ so that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$.
@@
[…]
@@
Figure 3: Two examples of the degree-reduction construction for a $g$ graph, respecting a rotation system, local to a vertex. The vertex is replaced with $\Delta$ vertices, at least one vertex for each incident edge, and these new vertices are connected with a near-perfect binary tree. In the left example, the displayed vertex is of maximum degree in its graph, while in the right example it is not (resulting in some degree-111 vertices).
@@
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated $g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $O(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
@@
Let $H_{1}$ be the graph that adds $\Delta$ vertices around each vertex to $G$, and ensure that between any two edges in the rotation system around a vertex in $G$, at least one such edge is added. Observe that $H_{1}$ is a $g$ graph with $(\Delta+1)n$ vertices and maximum degree $2\Delta$. Observe also that $G$ is a uniform shallow minor of $H_{1}$ of depth 222, and hence Lemma 3.11 shows that $\gamma^{(1)}(G)\lesssim(\Delta+1)\gamma^{(1)}(H_{1})$.
Now, we will add edges to $H_{1}$. In particular, we will construct a sequence of graphs that add edges to $H_{1}$, but never add vertices, and hence $\gamma^{(1)}$ will only increase for this sequence of graphs. Specifically, within each face of $H_{1}$, we create a cycle of all the newly added vertices in order around the face. The result is a graph we call $H_{2}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{3,2\Delta\}$) that has no adjacent pair of faces both with size >4>4> 4. Moreover, any vertex incident to a face of size >4>4> 4 (i.e., those in the created cycles) is incident exactly two other faces, and has degree exactly 333. For each face of exactly size 444 in $H_{2}$, triangulate it arbitrarily to get another new graph $H_{3}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{5,4\Delta\}$). See Figure 4 for examples of each of these graphs around a particular face. It’s worth noting that for non-cellular embeddings of higher genus graphs, it is possible that a non-simple face may include the same edge twice in the same “direction” during a traversal around the face (as opposed to the example in the figure, where the doubly-included edge is included in opposite directions). Fortunately, this does not cause issues for the construction described here either. It is also worth noting that these steps are necessary even if all the faces are simple: Specifically, we would like to handle the remaining non-triangular faces independently, and the separation ensures that we are not able to later produce parallel edges (see [KB97, Figure 6.3(b)] for a demonstration of the issues of adjacent simple faces).
It remains to triangulate the remaining non-triangular faces of $H_{3}$, none of which are adjacent, and all of which are simple. This can be accomplished by adding at most 222 edges per vertex. In particular, Kant and Bodlaender [KB92, KB97] gave an algorithm called “zig-zagging” that adds edges to a single face using a sequence of ear-cuts, triangulating the face. This algorithm applies in the case of faces on any rotation system, not just a rotation system on a planar graph. Call the final graph $H$, which is a triangulated $g$ graph with maximum degree $\max\{7,4\Delta\}$. See the last diagram in Figure 4 for a demonstration of this algorithm.
@@
For the remaining lemmas, we need only consider triangulated $g$ graphs.
A $g$ graph naturally embeds into a $g$ surface. Moreover, a $g$ graph in fact has a circle packing on a $g$ surface. Both the surface and packing is unique up to a certain set of transformations [HS93]. However, to bound $\gamma^{(3)}$ in a similar manner to Theorem 3.5, we would need a circle packing on the unit sphere $S^{2}$, likely of ply $O(g)$. Unfortunately, as Kelner [Kel04, Kel06b, Kel06a] pointed out, such circle packings need-not exist. Instead, in order to bound $\lambda_{2}$, Kelner made use of two steps: First, a specific subdivision procedure is applied to the graph, and it is shown that a bound for the graph obtained by a sequence of subdivision procedures also induces a bound on the initial graph. Second, by applying this procedure a sufficient number of times, it is shown that a continuous analogue of the circle packing theory can be used to obtain a bound. We will apply analogues of each of these steps for $\gamma^{(3)}$ (and $\gamma^{(1)}$, which differs only by a constant factor of at most 333) instead of $\lambda_{2}$.
@@
Let $G$ be a graph of genus $g$. For each $k$, let $F^{(k)}$ denote a $g$ surface on which $G^{(k)}$ admits a circle packing. For each vertex $v\in G^{(k)}$, let $C_{v}$ denote the disk in the packing corresponding to $v$, and let $r_{v}$ denote its radius. Let $A(C)$ and $D(C)$ denote the area (measure) and diameter (longest geodesic path between a point pair), respectively, of a compact and connected region $C$ either in $F^{(k)}$ or $S^{2}$. Then there is a sequence $f^{(k)}:F^{(k)}\to S^{2}$ of analytic maps so that, for any $\varepsilon>0$, there exists a threshold $N$ such that for any $k\geq N$ a partition of $S^{(k)}$ into $S^{(k)}_{1}\cup S^{(k)}_{2}$ exists with the following properties:
@@
[…]
@@
4 Bounds via 111-dimensional Extremal $L_{p}$-Embedded Spread
In this section, we will study a quantity we call the $d$-dimensional extremal $L_{2}$-embedded spread (with a focus on the 111-dimensional case). In more generality:
@@ 4 Bounds via 111-dimensional Extremal $L_{p}$-Embedded Spread
Apply Theorem 2.15 to get a non-expansive embedding $f^{\prime}:V\to\mathbb{R}$ from $d_{G}^{\omega}$ to $\mathbb{R}$ with average 222-distortion at most $C_{2}\cdot 2\cdot\left[\overline{\alpha}(G)\right]^{2}$ for a universal constant $C_{2}$. Using $y(v):=\omega(v)^{2}$ and $f(v):=f^{\prime}(v)/2$, we claim the constraints of $\overline{Q}_{2}^{(1)}(G)$ are satisfied: For the first constraint, $||y||_{1}=\sum_{v\in V}\omega(v)^{2}=||\omega||_{2}^{2}\leq 1$. For the remaining constraints,
@@
and so $y,f$ forms a feasible solution to $\overline{Q}_{2}^{(1)}(G)$. Finally, it follows from the 222-distortion bound that
@@
The proof is essentially the same as that of Lemma 4.5, except that a non-expansive embedding with bounded 111-distortion is used. ∎
@@ 6 A High-Level View of Congestion-based Bounds
[…]
@@
$\lambda_{2}^{*}(G)\lesssim(\log g)^{2}\cdot\frac{g}{n}$$\gamma^{(n)}(G)\lesssim(\log g)^{2}\cdot\frac{g}{n}$$\gamma^{(1)}(G)\lesssim(\log g)^{4}\cdot\frac{g}{n}$$\psi(G)\lesssim\frac{(\log g)^{2}\sqrt{g}}{\sqrt{n}}$trong duality Metric embedding/dimension reduction (Theorem 5.2) 1 11-dimensional Cheeger inequality for $\psi$ (Theorem 1.5)
@@ 7 Conclusion and Future Work
  • Can these methods be leveraged to efficiently obtain near-optimal balanced separators? The methods of Orecchia and Vishnoi [OV11], and Orecchia, Sachdeva, and Vishnoi [OSV12] seem to be promosing avenues for this possibility.
@@
Experimentally testing the practicality of the method would first require a complete implementation, so we outline the simplest-known methods for each step: Boyd, Diaconis and Xiao [BDX04] provided a practical and simple subgradient method for computing $\lambda_{2}^{*}$ that can quite easily be implemented with GPU-acceleration in a modern library. Unfortunately, their method as-stated is a primal method only, and does not obtain a dual solution, so a slightly modified form would be necessary. The simplest dimension-reduction technique is that of Kwok, Lau, and Tung [KLT22], stated in Theorem 1.6. Unfortunately, the best-known bounds for $\alpha(G)$ (particularly for forbidden-minor graphs) are not currently as algorithmically simple or practical, nor is their application in Theorem 5.2 and Theorem 5.3. Finally, the simplest cut (rounding) algorithm is also likely that of Kwok, Lau, and Tung [KLT22, Section 3.2.3], although its constants could certainly use some improvement.
8 Acknowledgements
Thank you to Therese Biedl for pointing out that not every embedded graph in the torus can be simply triangulated, which led to the $K_{5}$ example discussed at the start of Section 3.2.